a bag contains red and blue marbles. you pick two marbles at random without replacing the first one. The probability of drawing a blue and then a red is 4/15. The probability that your second marble is red if your first marble is blue is 2/3. Find the probability that the first marble is blue.
Answer by stanbon(75887)

A bag contains red and blue marbles. You pick two marbles at random without replacing the first one. The probability of drawing a blue and then a red is 4/15. The probability that your second marble is red if your first marble is blue is 2/3. Find the probability that the first marble is blue.
====
P(red | blue) = P(blue and red)/P(blue) = 2/3
-----
P(blue and red) = 4/15
===========
So, P(blue) = (4/15)/(2/3) = (4/15)*(3/2) = 2/5
=====================
Cheers,
Stan H.
=====================